# Learning the Covariance of Treatment Effects Across Many Weak Experiments

### v1

@@ 3.3 Nonparametric IV Model
$S$ $=\mu_{S}(T)+\pi_{S}(T)A+\eta$ (25)
$Y$ $=\mu_{Y}(T)+h(S)+\epsilon,$ (26)
where $E[\epsilon\mid A,T]=0$. We now show that, under some assumptions, $\theta_{1}$ identifies a functional of $h$.
@@
$\left\lVert\tau_{S}(t)\right\rVert_{2}\leq\epsilon$ for every $t\in[K]$ for some $\epsilon>0$.
@@
$\theta_{1}(\Lambda_{K})=$ $\left(\sum_{t=1}^{K}\pi_{S}(t)\pi_{S}(t)^{\top}\right)^{-1}\sum_{t=1}^{K}\pi_{S}(t)\pi_{S}(t)^{\top}E[\nabla h(S(t,0))]$ (27)
$+O(M\epsilon),$ (28)
@@
$\theta_{1}(\Lambda_{K})$ (29)
$=$ $\left(\sum_{t=1}^{K}\tau_{S}(t)\tau_{S}(t)^{\top}\right)^{-1}\sum_{t=1}^{K}\pi_{S}(t)E[h(S(t,0)+\pi_{S}(t))-h(S(t,0))]$ (30)
$=$ $\left(\sum_{t=1}^{K}\pi_{S}(t)\pi_{S}(t)^{\top}\right)^{-1}\sum_{t=1}^{K}\pi_{S}(t)\pi_{S}(t)^{\top}E[\nabla h(S(t,0))]$ (31)
$+\left(\sum_{t=1}^{K}\pi_{S}(t)\pi_{S}(t)^{\top}\right)^{-1}\sum_{t=1}^{K}\pi_{S}(t)\pi_{S}(t)^{\top}E[\nabla^{2}h(\widetilde{S}(t,0))\tau_{S}(t)],$ (32)
@@
4 Estimating Treatment Effect Covariances
@@
4.1 A Naive Estimator
A naive way to estimate the covariance matrix $\Lambda_{K}$ of true treatment effects is just to form the empirical covariance matrix $\hat{\Sigma}_{K}$ of the estimated ATEs $\widehat{\tau}_{S}(t)$ and $\widehat{\tau}_{Y}(t)$:
$\widehat{\Sigma}_{K}=\frac{1}{K}\sum_{t=1}^{K}\widehat{\tau}(t)\widehat{\tau}(t)^{\top}-\left(\frac{1}{K}\sum_{t=1}^{K}\widehat{\tau}(t)\right)\left(\frac{1}{K}\sum_{t=1}^{K}\widehat{\tau}(t)\right)^{\top},$ (33)
@@ 4.1 A Naive Estimator
$E\widehat{\Sigma}_{K}=\Lambda_{K}+\frac{4}{n}\bar{\Omega}_{K},$ (34)
@@
$\bar{\Omega}_{K}=$ $\frac{1}{K}\sum_{t=1}^{K}\frac{\Omega_{t,1}+\Omega_{t,0}}{2}$ (35)
@@
In the setting of industrial (especially digital) experimentation, experiment ATEs rarely exhibit large signal-to-noise ratio, meaning that $4\bar{\Omega}_{K}/n$ is often not negligible in front of $\Lambda_{K}$ Cunningham and Kim (2020); Peysakhovich and Eckles (2018). As a result, under our first or second structural model, the estimate of $\beta$ based on the empirical covariance matrix, $\widehat{\beta}=\widehat{\Sigma}_{SS,K}^{-1}\widehat{\Sigma}_{SY,K}$, is biased and is inconsistent for $\beta$ even as we gather more data in the form of additional experiments (that is, as $K\to\infty$). Since the two-stage least squares estimator under a categorical instrument equals OLS on the group means, it holds that $\widehat{\beta}$ is 2SLS with
@@
That $\widehat{\beta}$ is biased under non-negligible noise is the well-known fact that 2SLS is inconsistent under weak instruments Angrist et al. (1999). This suggests that ideas from the weak instrumental variable literature could help in better estimating $\Lambda_{K}$. In the next sections, we will consider three different methods inspired by weak IV estimators, respectively, (1) the Jackknife Instrumental Variable estimator (JIVE), (2) the Limited Information Maximum Likelihood (LIML) estimator, and (3) the general form of $k$-class estimators.
4.2 Covariance Matrix of Jackknifed Treatment Effects
We propose an estimator of $\Lambda_{K}$ inspired by the JIVE estimator Angrist et al. (1999). Introduce the transformed vector $\widetilde{D}=2(2A-1)D$. Observe that $E[\widetilde{D}\mid T=t]=\tau(t)=E[D\mid T=t,A=1]-E[E\mid T=t,A=0]$.
Figure 1: How Measurement Error Distorts Treatment Effect Covariances.
Figure 2: Effect of Rotating Estimated Treatment Effects by $\Omega^{-1/2}$.
@@ 4.2 Covariance Matrix of Jackknifed Treatment Effects
$\widehat{\tau}(t)=\frac{1}{n}\sum_{i:T_{i}=t}\widetilde{D}_{i}\qquad\text{and}\qquad\widehat{\tau}_{-i}(t)=\frac{1}{n-1}\sum_{\begin{subarray}{c}j\neq i\\ T_{j}=t\end{subarray}}\widetilde{D}_{j}$ (38)
@@
$\widehat{\Lambda}_{2}(t)=\frac{1}{n}\sum_{i:T_{i}=t}\widehat{\tau}_{-i}(t)\widetilde{D}_{i}^{\top}$ (39)
@@
$\widehat{\Lambda}_{K}^{\mathrm{JK}}=\frac{1}{K}\sum_{t=1}^{K}\widehat{\Lambda}_{2}(t)-\left(\frac{1}{K}\sum_{t=1}^{K}\widehat{\tau}(t)\right)\left(\frac{1}{K}\sum_{t=1}^{K}\widehat{\tau}(t)\right)^{\top}$ (40)
be the Jackknifed covariance matrix. The Jackknife construction ensures that only the common source of variation between units in the same experiments is captured: the variation due to the treatment assignment, as opposed to unit-level noise. It is immediate that the Jackknifed second-order moment matrix is an unbiased estimate of the second-order moment matrix of the true treatment effects. It can also readily be checked that under mild conditions the second order term has bias $O(N^{-1})$; that is, the bias scales with the total number of units in all experiments as opposed to the number of observations per experiment. We state this formally in the following proposition:
@@
As a corollary of the above fact, if they admit a probability limit, the plug-in estimator $\theta_{i}(\widehat{\Lambda}_{K}^{\mathrm{JK}})$, for either $i=1$ or $i=(2,\Psi)$, is a consistent estimate of the parameter $\theta_{i}(\Lambda_{K})$ as the number of experiments grows, $K\to\infty$, even as the size of each experiment remains fixed. The former is equivalent to the JIVE estimator, while the latter is equivalent to the estimator proposed in Hahn et al. (2004).
Note that the Jackknifed covariance matrix estimator does not require homoskedasticity of the unit-level noise – that is, it does not require $\Omega_{a,t}=\Omega$ for all $t,a$ – for consistency. However, its downside is that it requires us to Jackknife the unit-level data in every experiment, which can be computationally prohibitive when $n$ and $K$ are large. In the next two sections, we will on the contrary see how we can leverage homoskedasticity when it turns out to be a realistic assumption.
4.3 Extracting Features of Treatment Effect Covariances by Isotropizing Noise
We will assume in the next two sections that there exists $\Omega$ such that $\Omega_{t,a}=\Omega$ for every $t,a$. We will further assume that we know $\Omega$ to a high relative precision. These are often realistic assumptions in digital experiments: while treatment effects are small, correlations across metrics tend to be (1) stable across experiments, and (2) non-negligible, and thus easy to estimate with a high signal-to-noise ratio by leveraging populations across multiple experiments (e.g., the entire user base). Furthermore, considering metrics that are sufficiently non-redundant ensures that $\Omega$ is well-conditioned.
@@ 4.3 Extracting Features of Treatment Effect Covariances by Isotropizing Noise
$E\widehat{\Sigma}_{K}=\Lambda_{K}+\frac{4}{n}\Omega.$ (41)
@@
$E\Omega^{-1/2}\widehat{\Sigma}_{K}\Omega^{-1/2}=\Omega^{-1/2}\Lambda_{K}\Omega^{-1/2}+\frac{4}{n}I_{G}.$ (42)
@@
As our notation suggests, $\widehat{\theta}^{LIMLK}$ is equal to the LIMLK (LIML with Known noise covariance matrix) estimate with dependent variable $\widetilde{Y}=2(2A-1)Y$, endogenous predictors $\widetilde{S}=2(2A-1)S$, and instrument $T$. (This can be observed directly from the definition of LIMLK Anderson et al. (2009).)
As the smallest eigenvector of the covariance matrix of observations is the statistical target of Total Least Squares (TLS), the procedure we just described, and thus LIMLK, can be implemented as: (1) transform the observations by applying $\Omega^{-1/2}$, (2) run TLS in the transformed space, and (3) transform the smallest eigenvector outputted by TLS by applying $\Omega^{-1/2}$. We illustrate this procedure visually in Section 5.
@@
4.4 Estimating Treatment Effect Covariances by Subtracting Noise
In the previous subsection, we have only leveraged the direction of the known $\Omega$ but not its scale, allowing us to make a connection to LIMLK. (Note that the above procedure would give the same result if we used $\rho\Omega$ for $\rho>0$ instead of $\Omega$.) Under known $\Omega$, there is a perhaps more direct way to obtain an unbiased estimate of $\Lambda_{K}$: subtract $(4/n)\Omega$ from $\widehat{\Sigma}_{K}$. Formally, letting $\widehat{\Lambda}_{K}^{\mathrm{TC}}=\widehat{\Sigma}_{K}-(4/n)\Omega$ (where TC stands for Total Covariance), we have that
$E\widehat{\Lambda}_{K}^{\mathrm{TC}}=\Lambda_{K},$ (43)
and therefore, under the existence of the appropriate probability limits, the plug-in estimate $\theta_{i}(\widehat{\Lambda}_{K}^{\mathrm{TC}})$, for either $i=1$ or $i=(2,\Psi)$, is a consistent estimate of $\theta_{i}(\Lambda_{K})$. In particular, when either $S$ fully mediates the effect of $A$ on $Y$ (no direct effects) or when direct effects follow the INSIDE assumption, we can consistently estimate the structural parameter $\beta$ with $\theta_{1}(\widehat{\Lambda}_{K}^{\mathrm{TC}})$.
@@ 4.4 Estimating Treatment Effect Covariances by Subtracting Noise
$\theta_{1}(\widehat{\Lambda}_{K}^{\mathrm{TC}})=\frac{(\widetilde{S}^{0})^{\top}(I-(1+4/n)M_{T})\widetilde{Y}^{0}}{(\widetilde{S}^{0})^{\top}(I-(1+4/n)M_{T})\widetilde{S}^{0}},$ (44)
@@
with $M_{T}=I-P_{T}$, $P_{T}=\widetilde{T}(\widetilde{T}^{\top}\widetilde{T})^{-1}\widetilde{T}^{\top}$, and $\widetilde{T}$ is the $N\times K$ matrix of one-hot encodings of experiment membership $T$ (that is $\widetilde{T}_{i,t}=\bm{1}\{T_{i}=t\}$. One might recognize from (44) that $\theta_{1}(\widehat{\Lambda}_{K}^{\mathrm{TC}})$ is a so-called $k$-class IV estimator, with $k=1+4/n$.
5 Simulation Study
@@
To provide insight into our statistical setup and the performance of our estimators, we conduct a simulation study. The parameters of our simulations are chosen to reflect aspects of real-world data. The unit-level noise, represented by $\Omega$ in our statistical setup, is typically large relative to the variance-covariance matrix of treatment effects, represented by $\Lambda_{K}$, and is also potentially anti-correlated. For example, clicks may be positively correlated with conversions outside of any experiment, but treatments that increase “click-baitiness” can reduce conversions.
@@
Following this example, Figure 1 depicts a situation where treatment effects on a proxy metric $S$ are negatively correlated with treatment effects on a primary objective $Y$. Throughout the figure, the white arrow points in the direction of the covariance in true treatment effects. In the second panel, we plot the unit-level noise, which is positively correlated between $S$ and $Y$, represented by the black line. Lastly, in the third panel, we plot the sampling distribution of treatment effect estimates with a fixed sample size.
@@
As third panel of Figure 1 shows, the unit-level noise can overwhelm the treatment effect covariance when either the treatment effect covariance or the sample size is relatively small. As a result, naively estimating $\theta_{1}(\Lambda_{K})$ using the covariance matrix of the estimated treatment effects $\theta_{1}(\widehat{\Sigma}_{K})$ will be biased in the direction of $\theta_{1}(\Omega)$. In the absence of unit-level covariance (i.e., $\Omega=\omega I$), this bias is “merely” attenuation bias that preserves the direction of the relationship but biases estimates towards zero. In the presence of unit-level covariance, the estimated covariance can be biased in arbitrary directions, and this bias is worse for small experiments.
5.1 LIMLK as Total Least Squares

### v2

@@ 3.3. Nonparametric IV Model
(19) $S$ $=\mu_{S}(T)+\pi_{S}(T)A+\eta$
(20) $Y$ $=\mu_{Y}(T)+h(S)+\epsilon,$
where $E[\eta\mid A,T]=0$ and $E[\epsilon\mid A,T]=0$. We now show that, under some assumptions, $\theta_{1}$ identifies a functional of $h$.
@@
$\left\lVert\pi_{S}(t)\right\rVert_{\infty}\leq\epsilon$ for every $t\in[K]$ for some $\epsilon>0$.
@@
(21) $\theta_{1}(\Lambda_{K})=$ $\left(\sum_{t=1}^{K}\pi_{S}(t)\pi_{S}(t)^{\top}\right)^{-1}\sum_{t=1}^{K}\pi_{S}(t)\pi_{S}(t)^{\top}E[\nabla h(S(t,0))]$
(22) $+O(M\epsilon),$
@@
(23) $\theta_{1}(\Lambda_{K})$
(24) $=$ $\left(\sum_{t=1}^{K}\tau_{S}(t)\tau_{S}(t)^{\top}\right)^{-1}\sum_{t=1}^{K}\pi_{S}(t)E[h(S(t,0)+\pi_{S}(t))-h(S(t,0))]$
(25) $=$ $\left(\sum_{t=1}^{K}\pi_{S}(t)\pi_{S}(t)^{\top}\right)^{-1}\sum_{t=1}^{K}\pi_{S}(t)\pi_{S}(t)^{\top}E[\nabla h(S(t,0))]$
(26) $+\left(\sum_{t=1}^{K}\pi_{S}(t)\pi_{S}(t)^{\top}\right)^{-1}\sum_{t=1}^{K}\pi_{S}(t)\pi_{S}(t)^{\top}E[\nabla^{2}h(\widetilde{S}(t,0))\tau_{S}(t)],$
@@
4. Estimating Treatment Effect Covariances
@@
4.1. A Naive Estimator
A naive estimator for the covariance matrix $\Lambda_{K}$ of true treatment effects is simply the empirical covariance matrix $\hat{\Sigma}_{K}$ of the estimated ATEs $\widehat{\tau}_{S}(t)$ and $\widehat{\tau}_{Y}(t)$:
(27) $\widehat{\Sigma}_{K}=\frac{1}{K}\sum_{t=1}^{K}\widehat{\tau}(t)\widehat{\tau}(t)^{\top}-\left(\frac{1}{K}\sum_{t=1}^{K}\widehat{\tau}(t)\right)\left(\frac{1}{K}\sum_{t=1}^{K}\widehat{\tau}(t)\right)^{\top},$
@@ 4.1. A Naive Estimator
(28) $E\widehat{\Sigma}_{K}=\Lambda_{K}+\frac{4}{n}\bar{\Omega}_{K},$
@@
(29) $\bar{\Omega}_{K}=$ $\frac{1}{K}\sum_{t=1}^{K}\frac{\Omega_{t,1}+\Omega_{t,0}}{2}$
@@
In industrial (especially digital) experimentation, ATEs typically exhibit a low signal-to-noise ratio. This implies that the term $4\bar{\Omega}_{K}/n$ is often significant relative to $\Lambda_{K}$ (cunningham2020interpreting, ; peysakhovich2018learning, ). As a result, under our first or second structural model, the estimate of $\beta$ based on the empirica l covariance matrix, $\widehat{\beta}=\widehat{\Sigma}_{SS,K}^{-1}\widehat{\Sigma}_{SY,K}$, is biased and remains inconsistent for $\beta$ even as we increase the number of experiments (that is, as $K\to\infty$). Since the two-stage least squares estimator under a categorical instrument equals OLS on the group means, it holds that $\widehat{\beta}$ is 2SLS with
@@
The fact that $\widehat{\beta}$ is biased when noise is non-negligible reflects the well-established fact that 2SLS is inconsistent under weak instruments (angrist1999jackknife, ). This suggests that insights from the weak instrumental variable literature could improve the estimation of $\Lambda_{K}$. In the next sections, we will consider three different methods inspired by weak IV estimators: (1) the Jackknife Instrumental Variable estimator (JIVE), (2) the Limited Information Maximum Likelihood (LIML) estimator, and (3) the general form of $k$-class estimators.
4.2. Jackknifed Covariance Matrix of Treatment Effects
We propose an estimator of $\Lambda_{K}$ inspired by the JIVE estimator (angrist1999jackknife, ). Consider the transformed vector $\widetilde{D}=2(2A-1)D$. Observe that $E[\widetilde{D}\mid T=t]=\tau(t)=E[D\mid T=t,A=1]-E[E\mid T=t,A=0]$.
@@ 4.2. Jackknifed Covariance Matrix of Treatment Effects
(30) $\widehat{\tau}(t)=\frac{1}{n}\sum_{i:T_{i}=t}\widetilde{D}_{i}\qquad\text{and}\qquad\widehat{\tau}_{-i}(t)=\frac{1}{n-1}\sum_{\begin{subarray}{c}j\neq i\\ T_{j}=t\end{subarray}}\widetilde{D}_{j}$
@@
(31) $\widehat{\Lambda}_{2}(t)=\frac{1}{n}\sum_{i:T_{i}=t}\widehat{\tau}_{-i}(t)\widetilde{D}_{i}^{\top}$
@@
(32) $\widehat{\Lambda}_{K}^{\mathrm{JK}}=\frac{1}{K}\sum_{t=1}^{K}\widehat{\Lambda}_{2}(t)-\left(\frac{1}{K}\sum_{t=1}^{K}\widehat{\tau}(t)\right)\left(\frac{1}{K}\sum_{t=1}^{K}\widehat{\tau}(t)\right)^{\top}$
be the Jackknifed covariance matrix. The Jackknife construction ensures that only the common source of variation between units in the same experiments is captured: the variation due to the treatment assignment, as opposed to unit-level noise. It is immediate that the Jackknifed second-order moment matrix is an unbiased estimate of the second-order moment matrix of the true treatment effects. It can also readily be checked that under mild conditions the second order term has bias $O(N^{-1})$; in other words, the bias scales with the total number of units in all experiments as opposed to the number of observations per experiment. We state this formally in the following proposition:
@@
Proof.
As the second-order moment Jackknifed matrix $\allowbreak K^{-1}\allowbreak\sum_{t=1}^{K}\widehat{\Lambda}_{2}(t)$ is an unbiased estimate of $K^{-1}\sum_{t=1}\tau(t)\tau(t)^{\top}$, the bias reduces to that of the second term, which we rewrite as $N^{-1}Z_{N}Z_{N}^{\top}$, where $Z_{N}=N^{-1/2}\sum_{i=1}\widetilde{D}_{i}-E[\widetilde{D}_{i}]$. The (sequence of) random variables $Z_{N}$ can be controlled under various set of conditions. For example, suppose $0\leq m\preccurlyeq\Omega_{t,a}\preccurlyeq M<\infty$ for every $t,a$, and suppose that $K^{-1}\sum_{t}\Omega_{t,1}+\Omega_{t,0}\to\Omega$ in probability for some $\Omega$. Then, under a Lyapunov condition, a standard central limit theorem guarantees that $Z_{N}\rightsquigarrow Z$ where $Z\sim\mathcal{N}(0,\Omega)$. Then $E[(Z_{N}Z_{N}^{\top})]\to E[ZZ^{\top})]\preccurlyeq MI_{G}$, which implies the result. ∎
As a corollary, if it admits a probability limit, the plug-in estimator $\theta_{i}(\widehat{\Lambda}_{K}^{\mathrm{JK}})$ for either $i=1$ or $i=(2,\Psi)$ is a consistent estimator of the parameter $\theta_{i}(\Lambda_{K})$ as the number of experiments grows, $K\to\infty$, even as the size of each experiment remains fixed. The former is equivalent to the JIVE estimator, while the latter is equivalent to the estimator proposed in (hahn2004estimation, ).
Note that the Jackknifed covariance matrix estimator does not require homoskedasticity of the unit-level noise — that is, it does not require $\Omega_{a,t}=\Omega$ for all $t,a$ — for consistency. However, it does require us to Jackknife the unit-level data in every experiment, which can be computationally prohibitive when $n$ and $K$ are large. In the next two sections, we therefore consider how we can leverage homoskedasticity when it is a reasonable assumption.
4.3. Estimating Treatment Effect Covariances by Isotropizing Noise
In this section and the next, we assume a common noise covariance matrix $\Omega$ across all experiments and treatment arms, that is, $\Omega_{t,a}=\Omega$ for every $t,a$. We will further assume that we know $\Omega$ to a high relative precision. These are often reasonable assumptions in digital experiments: while treatment effects are small, correlations across metrics tend to be (1) stable across experiments, and (2) non-negligible, and thus easy to estimate with a high signal-to-noise ratio by leveraging populations across multiple experiments (e.g., the entire user base). Furthermore, considering metrics that are sufficiently nonredundant ensures that $\Omega$ is well-conditioned.
@@ 4.3. Estimating Treatment Effect Covariances by Isotropizing Noise
(33) $E\widehat{\Sigma}_{K}=\Lambda_{K}+\frac{4}{n}\Omega.$
@@
(34) $E\Omega^{-1/2}\widehat{\Sigma}_{K}\Omega^{-1/2}=\Omega^{-1/2}\Lambda_{K}\Omega^{-1/2}+\frac{4}{n}I_{G}.$
@@
As our notation suggests, $\widehat{\theta}^{LIMLK}$ is equal to the LIMLK (LIML with Known noise covariance matrix) estimate with dependent variable $\widetilde{Y}=2(2A-1)Y$, endogenous predictors $\widetilde{S}=2(2A-1)S$, and instrument $T$. (This can be observed directly from the definition of LIMLK (anderson2009limited, ).)
As the smallest eigenvector of the covariance matrix of observations is the statistical target of Total Least Squares (TLS), the procedure we just described, and thus LIMLK, can be implemented by: (1) transforming the observations by applying $\Omega^{-1/2}$, (2) running TLS in the transformed space, and (3) transforming the smallest eigenvector obtained from TLS by applying $\Omega^{-1/2}$. We illustrate this procedure visually in Section 5.
@@
4.4. Estimating Treatment Effect Covariances by Subtracting Noise
In the previous subsection, we leveraged only the direction of the known $\Omega$ but not its scale, allowing us to make a connection to LIMLK. (Note that the above procedure would give the same result if we used $\rho\Omega$ for $\rho>0$ instead of $\Omega$.) Under known $\Omega$, there is a perhaps more straightforward estimation procedure for $\Lambda_{K}$: subtract $(4/n)\Omega$ from $\widehat{\Sigma}_{K}$. Formally, letting $\widehat{\Lambda}_{K}^{\mathrm{TC}}=\widehat{\Sigma}_{K}-(4/n)\Omega$ (where TC stands for Total Covariance), we have that
(35) $E\widehat{\Lambda}_{K}^{\mathrm{TC}}=\Lambda_{K},$
and therefore, under the existence of the appropriate probability limits, the plug-in estimator $\theta_{i}(\widehat{\Lambda}_{K}^{\mathrm{TC}})$, for either $i=1$ or $i=(2,\Psi)$, provides a consistent estimator for $\theta_{i}(\Lambda_{K})$. In particular, when either $S$ fully mediates the effect of $A$ on $Y$ (no direct effects) or when direct effects follow the INSIDE assumption, we can consistently estimate the structural parameter $\beta$ with $\theta_{1}(\widehat{\Lambda}_{K}^{\mathrm{TC}})$.
@@ 4.4. Estimating Treatment Effect Covariances by Subtracting Noise
(36) $\theta_{1}(\widehat{\Lambda}_{K}^{\mathrm{TC}})=\frac{(\widetilde{S}^{0})^{\top}(I-(1+4/n)M_{T})\widetilde{Y}^{0}}{(\widetilde{S}^{0})^{\top}(I-(1+4/n)M_{T})\widetilde{S}^{0}},$
@@
with $M_{T}=I-P_{T}$, $P_{T}=\widetilde{T}(\widetilde{T}^{\top}\widetilde{T})^{-1}\widetilde{T}^{\top}$, and $\widetilde{T}$ is the $N\times K$ matrix of one-hot encodings of experiment membership $T$ (that is $\widetilde{T}_{i,t}=\bm{1}\{T_{i}=t\}$). One might recognize from (36) that $\theta_{1}(\widehat{\Lambda}_{K}^{\mathrm{TC}})$ is a so-called $k$-class IV estimator, with $k=1+4/n$.
5. Simulation Study
@@
To provide insight into our statistical setup and the performance of our estimators, we conduct a simulation study. The parameters of our simulations are chosen to reflect aspects of real-world data. The unit-level noise, denoted by $\Omega$ in our statistical setup, is typically large relative to the variance-covariance matrix of treatment effects, denoted by $\Lambda_{K}$, and is also potentially anti-correlated. For example, clicks may be positively correlated with conversions outside of any experiment, but treatments that increase “click-baitiness” can reduce conversions.
@@
Following this example, the top panel of Figure 1 depicts a situation where treatment effects on a proxy metric $S$ are negatively correlated with treatment effects on a primary objective $Y$. Throughout the figure, the white arrow points in the direction of the covariance in true treatment effects. In the middle panel, we plot the unit-level noise, which is positively correlated between $S$ and $Y$, represented by the black line. Lastly, in the bottom panel, we plot the sampling distribution of treatment effect estimates with a fixed sample size.
Figure 1. How Measurement Error Distorts Treatment Effect Covariances. This figure shows a hypothetical treatment effect covariance matrix, a unit-level sampling covariance matrix (“noise”), and the covariance of matrix of estimated treatment effects, which is a weighted combination of these.
@@
As bottom panel of Figure 1 shows, the unit-level noise can overwhelm the treatment effect covariance when either the treatment effect covariance or the sample size is relatively small. As a result, naively estimating $\theta_{1}(\Lambda_{K})$ using the covariance matrix of the estimated treatment effects $\theta_{1}(\widehat{\Sigma}_{K})$ will be biased in the direction of $\theta_{1}(\Omega)$. In the absence of unit-level covariance (i.e., $\Omega=\omega I$), this bias is “merely” attenuation bias that preserves the direction of the relationship but biases estimates towards zero. In the presence of unit-level covariance, the estimated covariance can be arbitrarily biased, and this bias is worse for small experiments (cunningham2020interpreting, ).
5.1. LIMLK as Total Least Squares
Figure 2. Effect of Transforming Estimated Treatment Effects by $\Omega^{-1/2}$. Transforming the covariance of estimated treatment effects renders the measurement error isotropic, so the TLS plim on this scatterplot coincides with the TLS target, whereas the OLS plim suffers from attenuation bias.